Daytime Company’s direct labor cost for February is as follows:

Actual direct labor hours 69,000

Total direct labor payroll $483,000

Efficiency variance $ 6,400 F

Rate variance $ 41,400 U

What were the standard direct labor hours?

A. 72,000

B. 71,400

C. 70,000

D. 69,000
Answer #1

Correct answer------------(C) 70,000

Working

For calculating standard hours we need to first calculate standard rate.

Labor Rate Variance
( Standard Rate - Actual Rate ) x Actual Labor Hours
( $                    6.40 - $                     7.00 ) x 69000
-41400
Variance $           41,400.00 Unfavourable-U

We know that rate variance is $41400 U. It means the standard rate is lower than actual rate.

The standard rate will be (41400/69000) = 0.60 lower than actual rate

The standard rate becomes (7-.60) = 6.40

.

Now we can calculate standard hours

Labour Efficiency Variance
( Standard Hours - Actual Hours ) x Standard Rate
( 70000 - 69000 ) x $              6.40
6400
Variance $             6,400.00 Favourable-F

Standard hours = Actual hours + (Labor efficiency variance / standard rate)

Standard hours = 69000+(6400/6.4)

Standard hours = 69000+1000

Standard hours = 70000 hours